OpenStudy (anonymous):

how do you solve this : 29-16+(-5) + ?

OpenStudy (aroub):

just add and subtract!

OpenStudy (aroub):

29-16-5=

OpenStudy (anonymous):

no

OpenStudy (aroub):

so whats the answer?!?!?

OpenStudy (anonymous):

i have a answer key but i have to show my work some how its 8

OpenStudy (aroub):

okay if u subtract 29-16-5 =8 / so my answer is correct! :)

OpenStudy (anonymous):

are you sure ?

OpenStudy (anonymous):

aroub is right! :)

OpenStudy (aroub):

yes im sure , thank you epaema!! :D

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

No problem! :)